HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

THE FIRST TORSIONALLY EXCITED STATE OF METHYLAMINE: ROTATIONAL SPECTRUM AND INTERSTELLAR DETECTION

P. GYAWALI, R. A. MOTIYENKO, UMR 8523 - PhLAM - Physique des Lasers

Atomes et Molecules, University of Lille, CNRS, F-59000 Lille, France; A. BELLOCHE, Millimiter und Submillimeter-Astronomie, Max-Planck-Institut fur Radioastronomie, Bonn, NRW, Germany; \underline{I. KLEINER} , Université Paris Cité and Univ Paris Est Créteil, CNRS, LISA, 75013, Paris France; V. ILYUSHIN, E. A. ALEKSEEV, Radiospectrometry Department, Institute of Radio Astronomy of NASU, Kharkov, Ukraine; I. GULACZYK, M. KREGLEWSKI, Faculty of

Chemistry, Adam Mickiewicz University, Poznan, Poland

Methylamine (CH_{3}NH_{2}) exhibits two large amplitude motions, methyl torsion and amine wagging, that complicate the spectral analysis especially in excited vibration states. To deal with the problem of fitting spectra in excited torsional states in the molecules like methylamine a new so-called ’hybrid’ Hamiltonian model was developed not far ago.^{1} Our current work aims to study experimentally and theoretically the terahertz rotational spectrum of methylamine to provide a reliable basis for the ISM detection of rotational transitions in v_{t} = 1 state. In its v_{t} = 0 state methylamine was first detected in the interstellar medium (ISM) toward Sgr B2 almost 50 years ago.^{2} The terahertz spectrum of methylamine was measured from 150 to 1520 GHz with the Lille fast scan spectrometer. Using a new hybrid Hamiltonian model, we were able to fit accurately the rotational spectrum of the v_{t}= 1 state of methylamine including the analysis of the nuclear quadrupole hyperfine structure. The results of this spectroscopic analysis allowed us to search for rotational transitions of methylamine in its first torsionally excited state toward the high-mass star forming region Sgr B2(N) that was the target of the imaging spectral line survey ReMoCA performed with the Atacama Large Millimeter/submillimeter Array (ALMA). We report the first interstellar detection of methylamine in its v_{t} = 1 state on the basis of this interferometric data set. Also, further steps in fitting the rotational transitions in the second torsional state of methylamine v_{t}= 2 will be shown.

This project has been supported by ANR Labex CaPPA through the PIA under Contract No. ANR-11-LABX-0005-01. This work was also supported by the Programme National de Physique et Chimie du Milieu Interstellaire (PCMI) of CNRS/INSU with INC/INP co-funded by CEA and CNES.

1Kleiner, I. and Hougen, J. T. 2015, J. Phys. Chem. A, 119, 10664

2Kaifu, et al. 1974, ApJ, 191, L135; Fourikis et al. 1974, ApJ, 191, L139 p-number: p104  Submitted on Tue, 13 Jun 2023 15:41:34 +0200